A particle P moves on a circle with centre O and radius b. At a certain instant the speed of P is "\upsilon" and its acceleration vector makes an angle "\alpha" with PO. Find the magnitude of the acceleration vector at this instant.


1
Expert's answer
2021-10-05T15:17:54-0400
"|a|=\dfrac{a_r}{\cos \alpha}=\dfrac{v^2}{b\cos \alpha}"